# Nonlinear Operator Learning Using Energy Minimization and MLPs

### v1

@@ Abstract
We develop and evaluate a method for learning solution operators to nonlinear problems governed by partial differential equations. The approach is based on a finite element discretization and aims at representing the solution operator by an MLP that takes latent variables as input. The latent variables will typically correspond to parameters in a parametrization of input data such as boundary conditions, coefficients, and right-hand sides. The loss function is most often an energy functional and we formulate efficient parallelizable training algorithms based on assembling the energy locally on each element. For large problems, the learning process can be made more efficient by using only a small fraction of randomly chosen elements in the mesh in each iteration. The approach is evaluated on several relevant test cases, where learning the solution operator turns out to be beneficial compared to classical numerical methods.
@@ 1 Introduction
In recent years, there has been a lot of interest in using machine learning methods in applied mathematics and scientific computing. In this work, the focus is on methods that approximates the solution operator to parameterized partial differential equations (PDEs). We emphasize that the aim is not to learn the solution to a single PDE problem, but rather to a family of the same type of PDE problems, by learning the mapping from a parameter space into the solution space.
There are several approaches to learning solution operators or solutions to PDEs in the literature. These can be divided into two main branches; data-driven or physics-informed. The data-driven methods are based on (typically pre-computed) data, such as e.g. Neural operators [6], Fourier neural operators [7], DeepONets [8], and the random feature model [9]. The physics-informed methods typically penalize the network to satisfy a PDE or other physical laws by designing an appropriate loss function, e.g. the Deep-Ritz method [4] and Physics-informed neural networks (PINNs) [11]. These methods do not require any data in general. Typically, the focus of physics-informed approaches has not been on learning parameterized problems, but mainly the solution to a single PDE problem.
@@
In this paper we develop a data-free physics-informed method for operator learning of parameterized PDEs. We construct a neural network that learns a family of solutions by utilizing a loss functions based on energy minimization that penalizes the network to satisfy the PDE. This means that the computationally costly step of generating training data is avoided. The energy minimization approach has earlier been considered in neural network contexts, e.g. [13] for finding low-dimensional models of physical motion and [12] for applications in soft tissue modeling. In this work we generalize these approaches and develop a method that can be applied to any (nonlinear) problem with a given energy functional. The problem is first discretized using a finite element method (FEM) and a neural network is trained to learn the corresponding approximate solution operator. The input to the neural network is typically parameters to the problem which defines the data on e.g. the boundary, the coefficients, or the right-hand side. The output of the network is the finite element degrees of freedom (DOFs), which for standard P1 elements considered here are the function values at the nodes of the finite element mesh.
@@
Furthermore, we provide a parallelized algorithm for computing the energy in each iteration, which is necessary for efficient training on GPUs. We also suggest to use random batches of elements in the mesh, which has potential to be useful in large problems with many degrees of freedom.
@@ 2 General setting and energy functionals
In this paper, we are interested in the solution operator to (2.1) depending on the parameter $\boldsymbol{p}$. Let $\mathcal{A}:D\to V$ denote this solution operator such that $u=\mathcal{A}(\boldsymbol{p})$. Here $\boldsymbol{p}$ is often referred to as the latent variable and $D$ the latent space. Furthermore, we assume that the latent variable $\boldsymbol{p}$ is a random variable with a given probability distribution $\mathcal{P}$ over $D$.
@@ 2.1 Finite element discretization
To be able to learn the solution operator we first need to discretize $V$. For this purpose, we introduce a finite element discretization of the solution space $V$. Let $\mathcal{T}_{h}$ denote a mesh based on a finite subdivision of $\Omega$ into closed and convex elements of maximal diameter $h$, i.e. $\mathrm{diam}(T)\leq h$ for $T\in\mathcal{T}_{h}$. Let $V_{h}\subseteq V$ be suitable finite element space based on the mesh $\mathcal{T}_{h}$. The corresponding approximation $u_{h}\in V_{h}$ is attained by minimizing the energy functional over the finite element space $V_{h}$
$u_{h}=\operatorname*{\arg\min}_{v\in V_{h}}E(v).$ (2.2)
@@
$\mathcal{A}_{h}:D\to V_{h},\quad\text{such that}\quad u_{h}=\mathcal{A}_{h}(\boldsymbol{p}).$ (2.3)
@@
We emphasize that the definition of $\mathcal{T}_{h}$ and $V_{h}$ are intentionally very general. The numerical examples in Section 4 will be based on first order Lagrangian (P1) finite elements on triangles and tetrahedrons. However, in the current framework, it is possible to use many other types of elements and also higher order degree polynomials.
@@ 3.1 Network architecture and loss function
We consider a simple feed-forward and fully connected network, often referred to as multilayer perceptron (MLP). The input layer is the latent variable $\boldsymbol{p}\in D$ and the output layer is the nodal values of a finite element function in $V_{h}$. The size of the input layer is $d$, since $\boldsymbol{p}\subseteq\mathbb{R}^{d}$, and the size of the output layer equals the number of free nodes $n$ in the mesh $\mathcal{T}_{h}$, i.e. nodes that do not correspond to a Dirichlet boundary condition.
@@
$\sigma(x)=\begin{cases}x,\quad&\text{if }x>0,\\ e^{x}-1,&\text{if }x\leq 0.\end{cases}$
Let $\theta$ denote the trainable weights in the network and let
$\mathcal{A}_{h,\theta}:D\to V_{h},\quad\text{such that}\quad u_{h,\theta}=\mathcal{A}_{h,\theta}(\boldsymbol{p})$
@@
To train the network we use the the following loss functions which is based on the energy
$\mathcal{L}(\theta):=\mathbb{E}_{\boldsymbol{p}\sim\mathcal{P}}\left[E(\mathcal{A}_{h,\theta}(\boldsymbol{p}))\right].$ (3.1)
@@
The network thus learns the operator by minimizing the expected value of the energy with respect to the parameter $\boldsymbol{p}$. Since minimizing the energy (for a fixed $\boldsymbol{p}$) is equivalent to solving the finite element problem, we expect the network to accurately approximate the (discretized) solution operator. This is further motivated in the next paragraph.
@@
The method is based on the idea of learning the finite element approximation. This turns out to be beneficial when studying the error of the method. First we observe that during training, the loss function $\mathcal{L}(\theta)=\mathbb{E}_{\boldsymbol{p}\sim\mathcal{P}}\left[E(\mathcal{A}_{h,\theta}(\boldsymbol{p}))\right]$ is indirectly minimized over a collection of functions in $V_{h}$ and $\mathbb{E}_{\boldsymbol{p}\sim\mathcal{P}}\left[E(\mathcal{A}_{h}(\boldsymbol{p}))\right]$ is the minimum over the set $V_{h}$. Hence, by assuming well-trainedness of the network we may assume that the error is small, i.e.
$\mathbb{E}_{\boldsymbol{p}\sim\mathcal{P}}\left[E(\mathcal{A}_{h,\theta}(\boldsymbol{p}))-E(\mathcal{A}_{h}(\boldsymbol{p}))\right]\leq\epsilon,$
for some $\epsilon>0$. Second, the error between the finite element solution and the exact solution is typically known from the rich finite element theory. We shall combine both of these ideas to obtain an approximation error estimate in Theorem 3.1. To quantify the error we assume that the energy takes the form
$E(v)=\frac{1}{2}(A\nabla v,\nabla v)_{L^{2}}-(f,v)_{L^{2}},$
@@
which is the case for a Poisson problem with coefficient $A$ and right-hand side $f$, see the numerical examples in Section 4 for further details. We also assume the that discretization is based on Lagrangian finite elements of order $p$. The error is measured in the following norm
$\|v\|^{2}_{a}:=a(v,v)=(A\nabla v,\nabla v)_{L^{2}},$
where we have the following identity.
Lemma 3.1 (An energy identity).
$\frac{1}{2}\|u_{h}-v\|_{a}^{2}=E(v)-E(u_{h})\quad\forall v\in V_{h}$
  • Proof.The right-hand side is
    $E(v)-E(u_{h})$ $=\bigg{(}\frac{1}{2}\|v\|_{a}^{2}-(f,v)_{L^{2}}\bigg{)}-\bigg{(}\frac{1}{2}\|u_{h}\|_{a}^{2}-(f,u_{h})_{L^{2}}\bigg{)}$
    $=\frac{1}{2}\|v\|_{a}^{2}-\frac{1}{2}\|u_{h}\|_{a}^{2}-(f,\underbrace{v-u_{h}}_{\in V_{h}})_{L^{2}}$
    $=\frac{1}{2}\|v\|_{a}^{2}-\frac{1}{2}\|u_{h}\|_{a}^{2}-a(u_{h},v-u_{h})$
    $=\frac{1}{2}\|v\|_{a}^{2}+\frac{1}{2}\|u_{h}\|_{a}^{2}-a(u_{h},v)$
    $=\frac{1}{2}\|u_{h}-v\|_{a}^{2}$
@@
Theorem 3.1 (An energy estimate for the approximation error).
There exists a positive constant such that
$\mathbb{E}_{\boldsymbol{p}\sim\mathcal{P}}\left[\|\mathcal{A}(\boldsymbol{p})-\mathcal{A}_{h,\theta}(\boldsymbol{p})\|_{a}^{2}\right]\lesssim h^{2p}\bigg{(}\mathbb{E}_{\boldsymbol{p}\sim\mathcal{P}}\left[\|D^{p+1}\mathcal{A}(\boldsymbol{p})\|^{2}\right]\bigg{)}+\epsilon$
  • Proof.Recall that $u=\mathcal{A}(\boldsymbol{p})$ and $u_{h,\theta}=\mathcal{A}_{h,\theta}(\boldsymbol{p})$. Hence, the argument to the expectation operator on left-hand side is bounded by
    $\|u-u_{h,\theta}\|_{a}^{2}$ $\lesssim\|u-u_{h}\|_{a}^{2}+\|u_{h}-u_{h,\theta}\|_{a}^{2}$
    $\lesssim h^{2p}\|D^{p+1}u\|^{2}+E(u_{h,\theta})-E(u_{h})$
    Here, we have started with an error split using ±$\pm u_{h}$, then applied a standard energy estimate for the finite element error and Lemma 3.1 for the learning error where we have used that $u_{h,\theta}\in V_{h}$. Taking the expected value of both sides preserves the inequality. Using linearity of the expected value and the well-trainedness assumption of the network gives the desired result. ∎

### v2

# Learning Nonlinear Finite Element Solution Operators using Multilayer Perceptrons and Energy Minimization

@@ Abstract
We develop and evaluate a method for learning solution operators to nonlinear problems governed by partial differential equations (PDEs). The approach is based on a finite element discretization and aims at representing the solution operator by a multilayer perceptron (MLP) that takes problem data variables as input and gives a prediction of the finite element solution as output. The variables will typically correspond to parameters in a parametrization of input data such as boundary conditions, coefficients, and right-hand sides. The output will be an approximation of the corresponding finite element solution, thus enabling support and enhancement by the standard finite element method (FEM) both theoretically and practically. The loss function is most often an energy functional and we formulate efficient parallelizable training algorithms based on assembling the energy locally on each element. For large problems, the learning process can be made more efficient by using only a small fraction of randomly chosen elements in the mesh in each iteration. The approach is evaluated on several relevant test cases, where learning the finite element solution operator turns out to be beneficial, both in its own right but also by combination with standard FEM theory and software.
@@ 1 Introduction
In recent years, there has been a lot of interest in using machine learning methods in applied mathematics and scientific computing, see for example [31]. In this work, we focus on methods that approximate the solution operator to parameterized partial differential equations (PDEs). We emphasize that the aim is not to learn the solution to a single PDE problem, but rather to a family of the same type of PDE problems, by learning the mapping from a low-dimensional parameter space into the solution space. As the solution space is typically infinite-dimensional we consider discrete spaces associated with standard numerical method for PDEs. We thus aim at learning well-known approximations instead of the exact solution. The idea is that the standard numerical method could support and enhance the machine learning one. Here we consider the finite element method (FEM) and present some examples of beneficial combination with the machine learning framework.
There are several approaches to learning solution operators or solutions to PDEs in the literature, see for example [19] where neural operators on a general integral form together with some concrete examples are presented. The different approaches can be divided into two main branches; data-driven or physics-informed. The data-driven methods are based on (typically pre-computed) data, such as, e.g., the neural operators in [21], Fourier neural operators [22], DeepONets [24], and the random feature model [26]. The physics-informed methods typically penalize the network to satisfy a PDE or other physical laws by designing an appropriate loss function, e.g., an energy-based loss function such as in the Deep-Ritz method [11] or a loss function based on the strong residual as for Physics-informed neural networks (PINNs) [30]. These methods do not require any data in general. Typically, the focus of physics-informed approaches has not been on learning parameterized problems, but mainly the solution to a single PDE problem, often with collocation type methods such as [11, 30, 7, 25, 36]. Recently however, works on learning parameterized PDEs seem to have gained more traction since they allow for a single network to predict the solutions to a class of related problems, see, e.g., [35, 13, 14, 17], or [10, 9] where PINNs is extended to parameterized settings, and [38] in which networks are trained to parameterize the Green function. Several parameterized-PDE approaches also take into account the variational formulation in some manner: [3] in a loss function related to the error in the norm induced by the variational formulation, [37] by basing the the loss function on the weak residual, and [29] instead in the network architecture. A work similar to the current is [17]. There, the map from a parameter space into a finite element space is also learned and some analysis is presented together with examples. However, there a field-to-field map is considered with convolutional neural networks (a natural choice for such a map) and both the analysis and examples are for linear problems. Here, we consider a tuple-to-field map with MLPs and both analysis and examples include nonlinearities. Another similar work is [13], where so called mesh-informed neural networks (MINNs) are used to map parameter tuples into a finite element space. There the loss function is based on the strong residual, not energy as here, also there is no analysis but several impressive examples are presented.
In this paper we consider a data-free physics-informed method for operator learning of parameterized PDEs. The method is actually a special case of the more general approach presented in [34] and it has already been applied to inverse problems in [8]. However, in the former, the method is barely noticeable in the more general framework presented there. In the latter, the focus lies on the application rather than on the method. Here, we therefore put full attention on the method itself, develop and present it in a general form along with some analysis and examples.
@@
We construct a neural network that learns a family of solutions by utilizing a loss function based on energy minimization that penalizes the network to satisfy the PDE. This means that the computationally costly step of generating training data is avoided. The energy minimization approach has earlier been considered in neural network contexts, e.g., [11, 33, 20, 12] and [34] for finding low-dimensional models of physical motion and [32] for applications in soft tissue modeling. In this work we consider a method that can be applied to any (nonlinear) problem with a given energy functional. The problem is first discretized using a finite element method and a multilayer perceptron (MLP) neural network is trained to learn the corresponding approximate solution operator. The input to the neural network is typically parameters to the problem which define the data on, e.g., the boundary, the coefficients, or the right-hand side. The output of the network is the finite element degrees of freedom (DOFs), which for, e.g., classical P1 elements are the function values at the nodes of the finite element mesh.
The general idea behind learning finite element solution operators is that regardless of machine learning method, a predicted PDE solution will be on a discrete form. It might therefore be reasonable to consider a discrete form related to a well-studied established method, rather than some seemingly arbitrary form specific to the learning approach. A motivation being that the established method might piggyback the machine learning one to enhance the final product. The established method we consider here is FEM due to its ubiquity, flexibilty, and substantial theory, but the general idea applies to any standard numerical method together with any machine learning framework. See Figure 1 for an illustration representing this general idea. The operator networks we consider thus map into a finite element space $V_{h}$ which allows this machine learning approach for PDEs to smoothly be combined with the powerful finite element method. This has a couple of advantages. Firstly, in analyzing the approximation error of the learned solution, tools from finite element theory may be incorporated to provide better error estimates. Secondly, existing finite element software may be used to improve the learned results. In this work we provide examples of both of these advantages.
@@
Furthermore, we provide a naturally parallelizable algorithm for computing the energy in each iteration, which is necessary for efficient training on GPUs. We also suggest to use random batches of elements in the mesh, which has potential to be useful in large problems with many degrees of freedom.
The paper is organized as follows. In Section 2 we present the general setting by introducing abstract spaces and energy functionals. In Section 3 we describe the neural network approach and define appropriate loss functions based on the energy. Finally, in Section 4 we present several numerical examples where operator learning turns out to be useful.
$u$
$u_{\theta}$
$\mathbb{R}^{n_{\text{output}}}$
$u$
$u_{h,\theta}$
$V_{h}$
$u_{h}$
Figure 1: Left: Simplified default approach to learning PDE-solution $u$. Right: Approach of instead learning the corresponding finite element solution $u_{h}$.
@@ 2 General setting and energy functionals
In this paper, we are interested in the solution operator to (2.1) depending on the parameter $\boldsymbol{p}$. Let $\mathcal{A}:D\to V$ denote this solution operator such that $u=\mathcal{A}(\boldsymbol{p})$. Here $\boldsymbol{p}$ is referred to as the problem data variable, or simply the parameter, and $D$ the parameter space. Furthermore, we assume that $\boldsymbol{p}$ is a random variable with a given probability distribution $\mathcal{P}$ over $D$.
@@ 2.1 Finite element discretization
To be able to learn the solution operator we first need to discretize $V$. For this purpose, we introduce a finite element discretization of the solution space $V$. Let $\mathcal{T}_{h}$ denote a mesh based on a finite subdivision of $\Omega$ into closed and convex elements of maximal diameter $h$, i.e., $\mathrm{diam}(T)\leq h$ for $T\in\mathcal{T}_{h}$. Let $V_{h}\subseteq V$ be suitable finite element space based on the mesh $\mathcal{T}_{h}$. The corresponding approximation $u_{h}\in V_{h}$ is attained by minimizing the energy functional over the finite element space $V_{h}$
$u_{h}=\operatorname*{\arg\min}_{v\in V_{h}}E(v)$ (2.2)
@@
$\mathcal{A}_{h}:D\to V_{h}\quad\text{such that}\quad u_{h}=\mathcal{A}_{h}(\boldsymbol{p})$ (2.3)
@@
We emphasize that the definition of $\mathcal{T}_{h}$ and $V_{h}$ are intentionally very general. The numerical examples in Section 4 will be based on first order Lagrangian (P1) finite elements on triangles. However, in the current framework, it is possible to use many other types of elements and also higher order degree polynomials.
@@ 3.1 Network architecture and loss function
We consider a simple feedforward and fully connected network, often referred to as multilayer perceptron. This choice of architecture is made since, to the best of our knowledge, no other work exists that combines the various components (MLP-approximated finite element solution operator, energy-based loss function, etc.) as is done here. We therefore think it is reasonable to first study and explore the basic standard architecture in this setting. Although basic, MLPs can posses good approximation capabilities as presented in the well-known work [15].
The input layer of the network is the problem data variable $\boldsymbol{p}\in D$ and the output layer is the nodal values of a finite element function in $V_{h}$. The size of the input layer is $d$, since $\boldsymbol{p}\subseteq\mathbb{R}^{d}$, and the size of the output layer equals the number of free nodes $n$ in the mesh $\mathcal{T}_{h}$, i.e., nodes that do not correspond to a Dirichlet boundary condition.
@@
$\sigma(x)=\begin{cases}x\quad&\text{if }x>0\\ e^{x}-1&\text{if }x\leq 0\end{cases}$ (3.1)
Let $\theta$ denote the trainable parameters (weights and biases) in the network and let
$\mathcal{A}_{h,\theta}:D\to V_{h}\quad\text{such that}\quad u_{h,\theta}=\mathcal{A}_{h,\theta}(\boldsymbol{p})$ (3.2)
@@
To train the network we use the the following energy-based loss function
$\mathcal{L}(\theta):=\mathbb{E}_{\boldsymbol{p}\sim\mathcal{P}}\left[E(\mathcal{A}_{h,\theta}(\boldsymbol{p}))\right]$ (3.3)
@@
The network thus learns the operator by minimizing the expected value of the energy with respect to the parameter $\boldsymbol{p}$. Since minimizing the energy (for a fixed $\boldsymbol{p}$) is equivalent to solving the finite element problem, we expect the network to accurately approximate the (discretized) solution operator. This is further motivated in Section 3.2. We remark that not all PDEs have a corresponding energy functional. For those that do not, one could instead use the final element residual during training, an example of weak residual minimization which is considered in [18, 5]. However, if available, the energy should be preferred to the residual. This is because the energy functional seems to be computationally cheaper to assemble and that it decomposes more easily into its local contributions, something that we exploit in Section 3.3.
3.2 Approximation error estimate
@@
The method is based on the idea of learning finite element solution operators. This turns out to be beneficial when studying the error of the method. First we observe that during training, the loss function $\mathcal{L}(\theta)=\mathbb{E}_{\boldsymbol{p}\sim\mathcal{P}}\left[E(\mathcal{A}_{h,\theta}(\boldsymbol{p}))\right]$ is indirectly minimized over a collection of functions in $V_{h}$ and $\mathbb{E}_{\boldsymbol{p}\sim\mathcal{P}}\left[E(\mathcal{A}_{h}(\boldsymbol{p}))\right]$ is the minimum over the set $V_{h}$. Hence, by assuming well-trainedness of the network we may assume that the error is small, i.e.,
$\mathbb{E}_{\boldsymbol{p}\sim\mathcal{P}}\left[E(\mathcal{A}_{h,\theta}(\boldsymbol{p}))-E(\mathcal{A}_{h}(\boldsymbol{p}))\right]\leq\epsilon$ (3.4)
[…]
$E(v)=\frac{1}{2}(A\nabla v,\nabla v)-(f,v)$ (3.5)
@@
which is the case for a Poisson problem with coefficient $A$ and right-hand side $f$. Here we consider both linear and nonlinear energies by letting $A$ depend on either the coordinates of the solution domain, like $v$, or depend on $v$ itself, respectively. We also assume the that discretization is based on Lagrangian finite elements of order $p$. The error is measured in the $H^{1}$-seminorm also informally called the $H_{0}^{1}$-norm
$|v|_{H^{1}}=\|v\|_{H_{0}^{1}}=\|\nabla v\|$ (3.6)
To obtain the approximation error estimate we will need the following results that connect the $H_{0}^{1}$-norm with the energy.
Lemma 3.1 (An identity for linear energy).
Let $E(v)=\frac{1}{2}a(v,v)-l(v)$, where $a$ is a bounded and coercive bilinear form on $V_{h}$ and $l$ is a bounded linear form on the same. Also let $a(u_{h},v)=l(v)$ for all $v\in V_{h}$, and $\|v\|_{a}^{2}:=a(v,v)$. Then
$\frac{1}{2}\|u_{h}-v\|_{a}^{2}=E(v)-E(u_{h})\quad\forall v\in V_{h}$ (3.7)
  • Proof.The left-hand side is
    $\frac{1}{2}\|u_{h}-v\|_{a}^{2}$ $=\frac{1}{2}\|v\|_{a}^{2}+\frac{1}{2}\|u_{h}\|_{a}^{2}-a(u_{h},v)$ (3.8)
    $=\frac{1}{2}\|v\|_{a}^{2}-\frac{1}{2}\|u_{h}\|_{a}^{2}+\|u_{h}\|_{a}^{2}-a(u_{h},v)$ (3.9)
    $=\frac{1}{2}\|v\|_{a}^{2}-\frac{1}{2}\|u_{h}\|_{a}^{2}+a(u_{h},\underbrace{u_{h}-v}_{\in V_{h}})$ (3.10)
    $=\frac{1}{2}\|v\|_{a}^{2}-\frac{1}{2}\|u_{h}\|_{a}^{2}+l(u_{h}-v)$ (3.11)
    $=\bigg{(}\frac{1}{2}\|v\|_{a}^{2}-l(v)\bigg{)}-\bigg{(}\frac{1}{2}\|u_{h}\|_{a}^{2}-l(u_{h})\bigg{)}$ (3.12)
    $=E(v)-E(u_{h})$ (3.13)
Under appropriate assumptions, the energy $E$ in (3.5) fulfills the requirements in Lemma 3.1 with $\|v\|_{a}^{2}=(A\nabla v,\nabla v)$. Thus applying Lemma 3.1 and noting that $\|v\|_{a}^{2}\geq\alpha\|\nabla v\|^{2}$ where $A\geq\alpha>0$, we have
$E(v)-E(u_{h})=\frac{1}{2}\|u_{h}-v\|_{a}^{2}\geq\frac{\alpha}{2}\|\nabla u_{h}-\nabla v\|^{2}$ (3.14)
We summarize this result in the following corollary.
Corollary 3.1 (An inequality for linear energy).
Let $E$ be defined by (3.5) with $A\geq\alpha>0$ and $f\in L^{2}(\Omega)$. Then there exists a positive constant such that
$\|\nabla u_{h}-\nabla v\|_{L^{2}(\Omega)}^{2}\lesssim E(v)-E(u_{h})\quad\forall v\in V_{h}$ (3.15)
Lemma 3.2 (An inequality for nonlinear energy).
Let $E(v)=\frac{1}{2}(A(v)\nabla v,\nabla v)=\frac{1}{2}\|\sqrt{A(v)}\nabla v\|^{2}$, i.e., $f=0$. Let $u_{h}\in V_{h}$ denote a minimizer to $E$, i.e., Gateaux derivative equal to zero for all $v\in V_{h}$. Assume that $A(v)\geq\alpha>0$ for all $v\in V_{h}$, and $A^{\prime}(u_{h})u_{h}\geq 0$. Then there exists a positive constant such that
$\|\nabla u_{h}-\nabla v\|_{L^{2}(\Omega)}^{2}\lesssim E(v)-E(u_{h})\quad\forall v\in V_{h}$ (3.16)
  • Proof.We start by scaling the left-hand side
    $\frac{\alpha}{2}\|\nabla u_{h}-\nabla v\|^{2}$ $=\frac{\alpha}{2}\|\nabla v\|^{2}+\frac{\alpha}{2}\|\nabla u_{h}\|^{2}-\alpha(\nabla u_{h},\nabla v)$ (3.17)
    $\leq\frac{1}{2}\|\sqrt{A(v)}\nabla v\|^{2}+\frac{1}{2}\|\sqrt{A(u_{h})}\nabla u_{h}\|^{2}-\alpha(\nabla u_{h},\nabla v)$ (3.18)
    $=E(v)+E(u_{h})-\alpha(\nabla u_{h},\nabla v)$ (3.19)
    $=E(v)-E(u_{h})+\underbrace{2E(u_{h})-\alpha(\nabla u_{h},\nabla v)}_{=R}$ (3.20)
    We consider the rest $R$ separately. For its treatment we use that one can write $A(v)=\alpha+a(v)$ for some function $a$ of $v$.
    $R$ $=2E(u_{h})-\alpha(\nabla u_{h},\nabla v)$ (3.21)
    $=(A(u_{h})\nabla u_{h},\nabla u_{h})-\alpha(\nabla u_{h},\nabla v)$ (3.22)
    $=(\alpha+a(u_{h})\nabla u_{h},\nabla u_{h})-\alpha(\nabla u_{h},\nabla v)$ (3.23)
    $=(\alpha\nabla u_{h},\nabla u_{h}-\nabla v)+(a(u_{h})\nabla u_{h},\nabla u_{h})$ (3.24)
    $\leq\alpha\|\nabla u_{h}\|\|\nabla u_{h}-\nabla v\|+(a(u_{h})\nabla u_{h},\nabla u_{h})$ (3.25)
    $\leq\frac{\alpha}{2\varepsilon}\|\nabla u_{h}\|^{2}+\frac{\varepsilon\alpha}{2}\|\nabla u_{h}-\nabla v\|^{2}+(a(u_{h})\nabla u_{h},\nabla u_{h})$ (3.26)
    We take $\varepsilon=1/2$ and kickback the second term. What remains is
    $\alpha\|\nabla u_{h}\|^{2}+(a(u_{h})\nabla u_{h},\nabla u_{h})=(A(u_{h})\nabla u_{h},\nabla u_{h})$ (3.27)
[…]
    =\;= 00 0 (3.29)
    Here we have used the assumption $A^{\prime}(u_{h})u_{h}\geq 0$ to get the inequality. The final step is noting that the obtained quantity is the left-hand side of the corresponding variational formulation (with test function $v=u_{h}$) which is equal to zero since $f=0$. Equivalently, the quantity is the Gateaux derivative of $E$ at $u_{h}$ (in the direction of $u_{h}$) which is equal to zero in any direction since $u_{h}$ is a minimizer to $E$. Summing up we have
    $\frac{\alpha}{4}\|\nabla u_{h}-\nabla v\|^{2}\leq E(v)-E(u_{h})$ (3.30)
    which is the desired inequality.
@@
Theorem 3.1 (An estimate for the approximation error).
For parameters $\boldsymbol{p}$ from a probability distribution $\mathcal{P}$, let $\mathcal{A}$ denote the solution operator to the continuous problem (2.1). Given a finite element discretization of this problem with mesh size $h$ and based on Lagrangian elements of polynomial degree $p$, let $\mathcal{A}_{h,\theta}$ denote the corresponding neural network approximation of the discrete solution operator $\mathcal{A}_{h}$. If we assume well-trainedness of the network with error bound $\epsilon>0$, i.e., (3.4), then there exists a positive constant such that
$\mathbb{E}_{\boldsymbol{p}\sim\mathcal{P}}\left[\|\nabla\mathcal{A}(\boldsymbol{p})-\nabla\mathcal{A}_{h,\theta}(\boldsymbol{p})\|_{L^{2}(\Omega)}^{2}\right]\lesssim h^{2p}\bigg{(}\mathbb{E}_{\boldsymbol{p}\sim\mathcal{P}}\left[\|D^{p+1}\mathcal{A}(\boldsymbol{p})\|_{L^{2}(\Omega)}^{2}\right]\bigg{)}+\epsilon$ (3.31)
  • Proof.Recall that $u=\mathcal{A}(\boldsymbol{p})$ and $u_{h,\theta}=\mathcal{A}_{h,\theta}(\boldsymbol{p})$. Hence, the argument to the expectation operator on left-hand side is
    $\|\nabla u-\nabla u_{h,\theta}\|^{2}$ $\lesssim\|\nabla u-\nabla u_{h}\|^{2}+\|\nabla u_{h}-\nabla u_{h,\theta}\|^{2}$ (3.32)
    $\lesssim h^{2p}\|D^{p+1}u\|^{2}+E(u_{h,\theta})-E(u_{h})$ (3.33)
    Here we have started with an error split using ±$\pm u_{h}$. Then we have applied a standard error estimate for the finite element error and a result like Corollary 3.1 or Lemma 3.2 for the learning error where we have used that $u_{h,\theta}\in V_{h}$. Taking the expected value of both sides preserves the inequality. Using linearity of the expected value and the well-trainedness assumption (3.4) of the network gives the desired result. ∎